Ling Li of Li​ Windows, Inc. is considering making a change in the material the firm uses for panes in its residential window line. The new material has a slight mirror attribute that assists in reflecting ultraviolet light and restricts the transmission of heat. The new material will raise the cost of a standard window by

​$3.423.42.

The current window is in the mature stage of the life​ cycle, and, with no​ modifications, Li has estimated that sales of the window line will be

250 comma 000250,000

units per year with a probability of

0.30.3

or

210 comma 000210,000

units per year with a probability of

0.70.7.

The standard net profit of a window unit is

​$5555.

With the new glass​ material, the selling price per unit can be increased by

​$55

​(but with the added cost per​ unit, the net profit will be

​(($5555plus+​$55​)minus−​$3.423.42​)equals=​$56.5856.58​).

​However, Li estimates that the demand for the newly designed window will be

220 comma 000220,000

units with a probability of

0.60.6​,

and there will be a

0.40.4

probability of sales of

180 comma 000180,000

units.

When using the existing​ material, the expected monetary value

​(EMV)equals=​$?????
Answer #1

Answer:

Existing material:

  • Expected sale = sum product of deamnd and their respective probability
  • Expected sale = 250000* 0.3 + 210000*0.7 = 222000 units
  • Expected profit = Expected sale* profit per unit = 222000*55 = $12,210,000 (Answer)